OpenStudy (anonymous):

hi i was hoping someone could help me write an expression for this problem, its pretty annoying

OpenStudy (anonymous):

just check out the attachment thanks

OpenStudy (anonymous):

The expression you're looking for is going to take the form y=mx+b. You need to find the value of m and b to complete the expression. The formula to find m is: \[m = \frac{ f(a) - f(b) }{ a-b }\] To find b you need to choose a set of values for x and y, either 10 shirts for $60 or 20 shirts for $100 and plug it into the equation y=mx+b and solve for b.

OpenStudy (anonymous):

i dont get it :I

OpenStudy (anonymous):

oh wait nevermind i got it, thanks!

OpenStudy (anonymous):

Let's start with m. They give us two quantities of shirts and how much they cost, so we need to plug that in to the formula I wrote above, which will give us: \[m= \frac{ $100-$60 }{ 20-10 }\] This will give you the slope of the line.

OpenStudy (anonymous):

You should get y=4x+20, or in other words, there is an initial cost of $20 and the price goes up $4 a shirt.

OpenStudy (anonymous):

yeah thats what i got